OpenStudy (anonymous):

simplify 15^-4 / 15^5

OpenStudy (anonymous):

\[15^{-4}\] the same as \[\frac{ 1 }{ 15^{4} }\]

OpenStudy (anonymous):

well since we are dividing and the base's are the same couldnt you just subtract the exponents and get the answer?

OpenStudy (anonymous):

yes, that's how to get the answer

OpenStudy (anonymous):

okay so -4 - 5 = -9 and there is no negative 9 theres only positive

OpenStudy (anonymous):

the result DOES have a negative exponent, it is possible - just like the problem gave us a term with negative exponent. negative power just means reciprocal of whatever base/exp you have

OpenStudy (anonymous):

-4-5 = -9, so the exponent is -9. it can be negative

OpenStudy (anonymous):

OK?